OpenStudy (anonymous):

A fraction has a denominator of 2 + 7i. What is multiplied with the fraction to rationalize the fraction?

OpenStudy (anonymous):

options: 2 + 7i 7 + 2i 2 – 7i 7 – 2i s

OpenStudy (anonymous):

You multiply by the conjugate, which is 2-7i.

OpenStudy (anonymous):

oh, so it's the first one?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I mean 3rd one

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Lets define : A/(2+7i) to get rid of the complex number in the denominator you must use the conjugate thats because of the difference of two square apply two complex numbers ( a+bi) (a-bi) = a^2 +b^2 so the answer will be 2-2i because it makes 4+49 =53 in the denominator

OpenStudy (anonymous):

thank you for clearing that up with me

OpenStudy (anonymous):

wait, don't you mean 2-7i?

OpenStudy (anonymous):